A customer purchased a care and financed $16,000. The customer borrowed part of the money from a bank charging 10% annual interest and the rest from a credit union at 8% annual interest. If the total interest for the year was $1390 how much borrowed from the bank and how much was borrowed from the credit uninon?
Found 2 solutions by stanbon, Earlsdon:
Answer by stanbon(57307)

A customer purchased a car and financed $16,000. The customer borrowed part of the money from a bank charging 10% annual interest and the rest from a credit union at 8% annual interest. If the total interest for the year was $1390 how much borrowed from the bank and how much was borrowed from the credit uninon?
----------------
Let amount invested at 10% be "x"; Interest on that is 0.10x dollars
------------
Amount invested at 8% is "16000-x"; Interest on that is 0.08(16000-x)=1280-0.08x dollars
-----------------
EQUATION:
interest + interest = $1390
0.10x + 1280 - 0.08x = 1390
0.02x = 110
x= $5500.00 (amt invested at 10%)
16000-x = 16000-5500 = $10500 (amt invested at 8%)
=================
Cheers,
Stan H.

Answer by Earlsdon(6287)

Let x = the amount borrowed from the bank at 10% interest.
Then $16,000-x = the amount borrowed from the credit union at 8% interest.
You can write the following equation, after changing the percentages to their eqivalent decimal amounts, for the total amount of interest paid:
x(0.1) + ($16000-x)(0.08) = $1,390 Simplify and solve for x.
0.1x + $1,280 - 0.08x = $1,390 Combine like-terms.
0.02x + $1.280 = $1,390 Subtract $1,280 from both sides.
0.02x = $110 Divide both sides by 0.02
x = $5,500 This is the amount borrowed from the bank at 10% interest.
$16,000 - x = $16,000 - $5,500 = $10,500 This is the amount borrowed from the credit union at 8% interest.